OpenStudy (anonymous):

i'm having a bit of trouble with U substitution like how do I do it exactly with a problem like X/1+X^4 dx

OpenStudy (anonymous):

Do you mean \[\int\frac{x}{1+x^4}dx?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then an ordinary algebraic substitution won't work here. Just a sec, let me try to work this out.

OpenStudy (anonymous):

okay then

OpenStudy (anonymous):

Alright, I think I have it. Make the following substitution: \[x=\sqrt{u}, \text{ so}\\ x^2=u, \text{ and}\\ x^4=u^2\] So, you have \[dx=\frac{1}{2\sqrt{u}}du\] Now the integral becomes \[\int\frac{\sqrt{u}}{1+u^2}\left(\frac{1}{2\sqrt{u}}du\right)\\ \frac{1}{2}\int\frac{1}{1+u^2}du\] Can you work with that one?

OpenStudy (anonymous):

i think so let me give it a try

OpenStudy (anonymous):

I have one question where did u get the 1/2\[1\2\sqrt{U}\]

OpenStudy (anonymous):

That's from the derivative of sqrt(u): \[\frac{d}{du}\sqrt{u}=\frac{d}{du}u^{\frac{1}{2}}=\frac{1}{2}u^{-\frac{1}{2}}=\frac{1}{2u^{\frac{1}{2}}}=\frac{1}{2\sqrt{u}}\]

OpenStudy (anonymous):

that makes it easier to see